OpenStudy (anonymous):

Martin has a compost pit in his garden. The pit is 9 feet long, 6 feet wide, and 2 feet deep. She plans to make it 1.5 feet deeper. What will be the increase in the volume of the pit? Answer 18 cubic feet 54 cubic feet 81 cubic feet 162 cubic feet

OpenStudy (anonymous):

81 cubic feet

OpenStudy (anonymous):

i got like 109.5 i know im rong

OpenStudy (anonymous):

the length and width of the pit remain the same, so we can consider it as a cuboid of l*b*h = 9*6*1.5 = 81

OpenStudy (anonymous):

thank you XD

OpenStudy (ash2326):

\[Volume= length\ \times breadth\ \times depth\] Earlier Volume \[Volume=9 \times 6 \times 2=108\ feet^3 \] Now depth is increased by 1.5 feet so \[New\ Volume=9 \times 6 \times 3.5=189 feet^3\] Increase ij Volume=189-108= 81 feet^3